# Bayesian Structure Learning in Undirected Gaussian Graphical Models: Literature Review with Empirical Comparison

### v2

@@ 3 Review of Methodology
Bayesian approaches in structure learning aim to uncover the posterior distribution $P(G|\mathbf{X})$ of the graph. For that, one could either evaluate the posterior (6) or the joint posterior (8). For both, the prior distribution $P(\mathbf{K}|G)$ is required. Early Bayesian structure learning focuses on proposing this prior. Ideally, this prior is conjugate, i.e., the prior and posterior follow the same distribution family. Dawid & Lauritzen (1993) define such prior and call it the Hyper Inverse-Wishart (HIW) prior. They show that, apart from conjugacy, this prior also satisfies the so-called weak and strong hyper Markov property. Using the HIW prior, Giudici & Green (1999) and Carvalho et al. (2007) design algorithms for Bayesian structure learning that output the posterior for small graphs. The HIW prior, however, is only defined on decomposable graphs, which form a small subset of the whole graph space. Roverato (2002) solves this by extending the HIW to general graphs, a distribution that Atay-Kayis & Massam (2005) later coin the G-Wishart distribution. Its density is given by
$P(\mathbf{K}|G)=\frac{1}{I_{G}(b,\mathbf{D})}|\mathbf{K}|^{\frac{b}{2}-1}\exp\{-\frac{1}{2}tr(\mathbf{K}\mathbf{D})\}I(\mathbf{K}\in\mathcal{P}_{G}),$ (10)
where $\mathcal{P}_{G}\subset\mathcal{K}$ denotes the set of symmetric positive definite matrices $\mathbf{K}$ that have $k_{ij}=0$ when $(i,j)\not\in G$. Here, $I(\mathbf{K}\in\mathcal{P}_{G})$ is an indicator function that is equal to one, if $\mathbf{K}\in\mathcal{P}_{G}$, and zero otherwise. The symmetric positive definite matrix $\mathbf{D}$ and the scalar $b>2$ are called the scale and shape parameters of the G-Wishart distribution. They are mostly set to $b=3$ and $\mathbf{D}=\mathbf{I}$. The normalizing constant
@@
Using the approximations for the normalizing constants in this ratio, one can approximate $P(\mathbf{X}|G)$. Jones et al. (2005) and Lenkoski & Dobra (2011) both design stochastic search algorithms that in each iteration favour a move to a graph with a high posterior probability $P(G|\mathbf{X})\propto P(\mathbf{X}|G)P(G)$. However, the approximation of the normalizing constants in (12) is unstable and its accuracy can be hard to access. Moreover, the approximations are computationally expensive, limiting structure learning to small graphs with less than 20 nodes.
That is why others focus on inference on the joint posterior (8). For this and for inference on the precision matrix in general, it is required to sample from the G-Wishart distribution (10), which is challenging due to the normalizing constant (11). This led to a string of articles on efficient sampling from the G-Wishart distribution, see Wang & Li (2012) for a review of those. However, sampling from the G-Wishart distribution remained computationally inefficient and unreliable.
@@ 3.1 MCMC-based Approaches
Most methods in Bayesian structure learning use MCMC sampling. MCMC is popular because of the variety of inferences it provides. Specifically, when using MCMC sampling, one can approximate the expected value of $h(G^{*})$ for any information function $h:\mathcal{G}\to\mathbb{R}$. For example, one can choose $h(G)$ equal to $1$ when an edge $e$ is in $G$, and 00 otherwise. This will give the posterior edge inclusion probabilities
@@
Similarly, with other information functions $h$, one can evaluate different posterior probabilities. For example, the probability that the true graph will have more than $10$ edges, the probability that the true graph is connected, the probability that the true graph contains a cycle, and so on. However, this flexibility comes at a cost: MCMC algorithms are computationally demanding because they iteratively explore large parameter spaces.
@@ 3.1.1 MCMC-based approaches on the joint space
Most of the MCMC approaches in Bayesian structure learning create a Markov Chain $((G^{(0)},\mathbf{K}^{(0)}),\dots,(G^{(S)},\mathbf{K}^{(S)}))$ over the joint space of graphs and precision matrices that converges to the joint posterior (8). Using Equation (14) these methods can perform inference on almost all characteristics of both the graph and the precision matrix. We call these methods joint MCMC methods. They come in three types: reversible jump (RJ), birth-death (BD), and spike-and-slab (SS). All joint MCMC methods are listed in Table 2.
@@
One can define the acceptance probabilities $\alpha(G^{\prime},G)$ such that the balance condition holds and the resulting MCMC chain converges to the joint posterior. The first three RJ methods to use this strategy are the RJ-Original (RJ-O) method by Dobra et al. (2011), the RJ-Wang Li (RJ-WL) method by Wang & Li (2012), and the RJ-Cheng Lenkoski (RJ-CL) method by Cheng & Lenkoski (2012). These methods form the foundation for all later reversible jump methods. Their approach to sample from the G-Wishart distribution, however, is not scalable to bigger graphs and lacks a theoretical guarantee of convergence. To address these issues, Lenkoski (2013) introduces an efficient exact sampler from the G-Wishart distribution. The resulting RJ-Double (RJ-D) method thereby becomes the first joint method to achieve detailed balance, and hence, theoretical convergence. Hinne et al. (2014) combine several existing techniques to improve the overall efficiency in the RJ-Double Continuous Bayes Factor (RJ-DCBF) method. Despite these improvements, RJ methods are still too computationally demanding for graphs with more than 50 variables. One of the reasons is that the acceptance probabilities contain a ratio of normalizing constants that is time-consuming to evaluate. Mohammadi et al. (2023a) tackles this by developing a closed-form approximation of the ratio used in the RJ - Approximation (RJ-A) method by Mohammadi & Wit (2019). Due to the approximation, the RJ-A method loses theoretical convergence but gains computational efficiency. The RJ-Weighted Proposal (RJ-WWA) method by Van den Boom et al. (2022a) combines the best of both worlds: a method that has both theoretical convergence and computational efficiency. Their improved strategy starts with two observations. First, at each iteration, all RJ algorithms propose a new graph that adds or removes an edge at random. This results in lots of graphs with low posterior probabilities that are proposed and subsequently rejected, thereby wasting computing power. The RJ-WWA algorithm is more likely to propose graphs with higher acceptance probabilities. The second observation is that the expensive calculation of the exact acceptance probability $\alpha(G^{\prime},G)$ wastes computing power, especially if the proposed graph $G^{\prime}$ gets rejected. Therefore, they introduce a pre-acceptance step, in which the acceptance probabilities are approximated using the computationally efficient approximation of Mohammadi et al. (2023a). Only when pre-accepted, the exact acceptance probabilities are calculated. This so-called delayed acceptance requires the expensive calculation of exact acceptance probabilities only for $G^{\prime}$ that are likely to be accepted, hence minimizing the lost computational effort. The most recent reversible jump method is the RJ-Unbiased (RJ-UB) method by Van den Boom et al. (2022b). The RJ-UB method introduces unbiased estimation to Bayesian structure learning by applying a technique of Glynn & Rhee (2014). It requires the construction of two Markov Chains $(G^{(0)},\cdots,G^{(S)})$ and $(\bar{G}^{(0)},\cdots,\bar{G}^{(S)})$ that meet at some time $\tau<\infty$, that is, $G^{(s)}=\bar{G}^{(s)}$ for all $s\geq\tau$. Now, the quantity
@@
is an unbiased estimator of $E(h(G^{*})|\mathbf{X})$. The main challenge in obtaining the unbiased estimator $\hat{h}_{q}$ is the construction of the two Markov chains that meet at a time $\tau$ and for which the balance conditions hold. To do so, Van den Boom et al. (2022b) use a combination of a sequential Monte Carlo (SMC) method and a particle-independent Metropolis-Hastings method. Van den Boom et al. (2022b) are not the first to use SMC techniques in Bayesian structure learning. Tan et al. (2017) also use an SMC sampler. Their method, however, is computationally cumbersome and comes with no theoretical guarantees.
The informed proposal and delayed acceptance mechanisms of the RJ-WWA method are one way of dealing with the low acceptance probabilities of the reversible jump approach. Mohammadi & Wit (2015) develop a different strategy to tackle this issue and create a new class of joint MCMC methods: Birth-Death (BD) methods. Their BD-Original (BD-O) method abolishes acceptance probabilities altogether while maintaining theoretical convergence. The BD-O method designs a continuous time (CT) Markov chain. This chain spends a continuous time $W(s)$ in each state $s$ before moving to the next state. The next state is determined by adding or removing an edge based on so-called birth-death rates that need to be recalculated at every iteration. Its pseudo-code is shown in Algorithm 2.
@@
for state $s=1,..,S$ do
       Calculate the birth-death rate $r_{e}$ for every edge $e\in G^{(s-1)}$;
@@
Output: Samples from the joint posterior (8)
Algorithm 2 Birth-death algorithm
By defining the birth-death rates correctly one can prove that the BD-O method meets the balance condition and therefore converges to the joint posterior. One can obtain the desired expected value using
@@
where $W$ is the sum of the waiting times of all states. Per iteration, the BD-O algorithm is computationally costly because it calculates the birth and death rates for all edges. However, its MCMC chain needs less iterations to converge because it moves to a new graph every iteration. Moreover, the calculation of the birth and death rates can be performed in parallel. Just like the acceptance ratio of the RJ methods, the birth and death rates of the BD-O method contain a ratio of normalizing constants. Using the approximation of this ratio by Mohammadi et al. (2023a), the BD-Approximation (BD-A) method achieves a significant reduction in computational cost. Due to this approximation the balance condition no longer holds and there is no theoretical convergence.
Despite all the mentioned improvements, RJ and BD methods still suffer from a high computational burden. This has two reasons: first, at every iteration, a new precision matrix needs to be sampled from the $G$-Wishart distribution, which remains, despite the direct sampler of Lenkoski (2013), computationally costly. Second, at every MCMC iteration, the graph changes by at most one edge. This leads to poor mixing and convergence of the MCMC chain, especially for large-scale graphs. Both problems were overcome by Wang (2015) with his Spike-and-Slab (SS) method. First, Wang (2015) defined a new prior on the precision matrix, called the spike-and-slab prior. Remember that the G-Wishart prior (10) sets elements $k_{ij}$ to zero when the corresponding element $(i,j)\not\in G$. The spike-and-slab (SS) prior samples these elements from a normal density with a variance $v_{0}$. The resulting elements are therefore not exactly zero, but, by choosing $v_{0}$ to be small, they are likely to be close to zero. The other elements are sampled from a normal density with a higher variance $v_{1}>v_{0}$. The elements of the precision matrix are still dependent, since the resulting matrix $\mathbf{K}$ needs to be positive definite. Moreover, like the G-Wishart prior, the SS prior contains an intractable normalizing constant. The SS prior allows for the design of an efficient block Gibbs sampler that updates the precision matrix column by column. Instead of adding or deleting a single edge (like the RJ or BD algorithms), the SS algorithm samples an entirely new graph at each iteration. The Markov chain’s mixing is therefore greatly improved compared to the RJ and BD methods. The pseudo-code of the SS algorithm is shown in Algorithm 3
@@
for state $s=1,..,S$ do
       Sample a new precision matrix $\mathbf{K}^{(s)}$ using block Gibbs sampling;
       Calculate the probabilities $l_{e}$ for every edge $e$;
@@
Output: Samples from the joint posterior (8).
Algorithm 3 Spike-and-slab algorithm
The probabilities $l_{e}$ in Algorithm 3 are calculated using an approximation. The balance condition therefore does not hold and the SS algorithm is not guaranteed to converge to the the joint posterior (8).
@@
Spike-and-slab Discrete SS Wang (2015)
Table 2: Overview of joint MCMC methods for Bayesian structure learning in GGMs with their corresponding prior on $\mathbf{K}$, type of Markov chain (discrete or continuous), name, and reference. The last column indicates whether an algorithm theoretically converges to the posterior distribution.
@@ 3.1.2 MCMC-based approaches on the graph space
Joint MCMC methods move over the joint space of graphs and precision matrices. An additional benefit of this strategy is that with the samples $(G^{(1)},\mathbf{K}^{(1)}),\dots,(G^{(S)},\mathbf{K}^{(S)})$, one can perform inference on the precision matrix. However, one is often just interested in retrieving the structure of the true graph. In that case, obtaining a new precision matrix at every iteration is a computational burden. The class of methods in this subsection overcomes this burden by only exploring the graph space. We also refer to these types of methods as G-methods.
Stingo & Marchetti (2014) introduce this strategy in their G-Stingo method. They start with the well-known result that if a graph $G_{d}$ is decomposable, the joint likelihood $P(\mathbf{X}|\mathbf{K},G_{d})$ can be written as the product of $p$ different regressions:
$P(\mathbf{X}|\mathbf{K},G_{d})$ =$\prod_{i=1}^{p}P(X_{i}|X_{-i},G_{d},\mathbf{K})=\prod_{i=1}^{p}\sum_{j\not=i}\beta_{ij}X_{j}+\epsilon_{i}.$ (17)
Note the connection with the frequentist regression approach in (3). They then observe that every general graph $G$ is a subgraph of a decomposable graph. Every graph $G$ can therefore be denoted by the pair $(G_{d},L)$, where $G_{d}$ is decomposable and $L$ is a subgraph of $G_{d}$. By putting priors on $\beta_{ij}$ and $\epsilon_{i}$, they then construct an MCMC algorithm that moves from $(G^{(s)},L^{(s)})$ to $(G^{(s+1)},L^{(s+1)})$ for $s=0,\dots,S-1$. This Markov chain approximates the posterior distribution $P(G,L|\mathbf{X})$, but the balance conditions do not hold. Therefore, the G-Stingo method does not have theoretical convergence.
Equation (17) only holds when the graph $G_{d}$ is decomposable. When $G$ is a general graph, this no longer holds. However, we can use the same strategy to find an approximation of the joint likelihood:
$P(\mathbf{X}|\mathbf{K},G)\approx\prod_{i=1}^{p}P(X_{i}|X_{-i},G,\mathbf{K}).$ (18)
The right-hand side of this expression is called the pseudo-likelihood. We can use it to approximate the marginal likelihood $P(\mathbf{X}|G)$:
@@
where $\hat{P}(\mathbf{X}|G)$ is called the marginal pseudo likelihood (MPL). Leppä-aho et al. (2017) introduces the concept of marginal pseudo likelihood to GGMs. They use the Wishart prior with $b=p$ and $D=\frac{1}{n}\mathbf{X}^{T}\mathbf{X}$. The Wishart prior is equivalent to the $G$-Wishart prior (10), where $G$ is the complete graph. Note that this prior uses the data $\mathbf{X}$. Such priors are called fractional priors. Leppä-aho et al. (2017) choose this prior because it leads to a closed-form expression of $\hat{P}(\mathbf{X}|G)$, and then design a hill-climbing algorithm.
The G-MPL-Reversible Jump (G-MPLRJ) method by Mohammadi et al. (2023b) introduce the closed-form expression of $\hat{P}(\mathbf{X}|G)$ in an MCMC algorithm. Their discrete time MCMC algorithm creates the chain $(G^{(0)},\cdots,G^{(S)})$. Its pseudo-code resembles Algorithm 1 with two differences. First, no precision matrix is sampled at every iteration. Second, the acceptance probabilities $\alpha(G^{\prime},G)$ are calculated using the marginal pseudo-likelihood, greatly improving the computational efficiency. The G-MPLRJ algorithm meets the balance conditions and therefore converges. However, the stationary distribution is not the true posterior $P(G|\mathbf{X})$, but instead $\hat{P}(G|\mathbf{X}):=\hat{P}(\mathbf{X}|G)P(G)/P(\mathbf{X})$. The same work also presents the G-MPL-Birth Death (G-MPLBD) method. Again using the approximation in (19), this method designs a continuous time MCMC similar to the BD-O algorithm. The resulting Markov chain also converges to $\hat{P}(G|\mathbf{X})$ instead of to the true posterior $P(G|\mathbf{X})$.

### v3

@@ 3 Review of Methodology
Bayesian approaches in structure learning aim to uncover the posterior distribution $P(G|\mathbf{X})$ of the graph. For that, one could either evaluate the posterior (6) or the joint posterior (8). For both, the prior distribution $P(\mathbf{K}|G)$ is required. Early Bayesian structure learning focuses on proposing this prior. Ideally, this prior is conjugate, i.e., the prior and posterior follow the same distribution family. Dawid & Lauritzen (1993) define such prior and call it the Hyper Inverse-Wishart (HIW) prior. Using the HIW prior, Giudici & Green (1999) and Carvalho et al. (2007) design algorithms for Bayesian structure learning that output the posterior for small graphs. The HIW prior, however, is only defined on decomposable graphs, which form a small subset of the whole graph space. Roverato (2002) solves this by extending the HIW to general graphs, a distribution that Atay-Kayis & Massam (2005) later coin the $G$-Wishart distribution. Its density is given by
$P(\mathbf{K}|G)=\frac{1}{I_{G}(b,\mathbf{D})}|\mathbf{K}|^{\frac{b}{2}-1}\exp\left\{-\frac{1}{2}tr(\mathbf{K}\mathbf{D})\right\}I(\mathbf{K}\in\mathcal{P}_{G}),$ (10)
where $\mathcal{P}_{G}\subset\mathcal{K}$ denotes the set of symmetric positive definite matrices $\mathbf{K}$ that have $k_{ij}=0$ when $(i,j)\not\in G$. Here, $I(\mathbf{K}\in\mathcal{P}_{G})$ is an indicator function that is equal to one, if $\mathbf{K}\in\mathcal{P}_{G}$, and zero otherwise. The symmetric positive definite matrix $\mathbf{D}$ and the scalar $b>2$ are called the scale and shape parameters of the $G$-Wishart distribution. They are mostly set to $b=3$ and $\mathbf{D}=\mathbf{I}$, where $\mathbf{I}$ is the p-dimensional identity matrix. The normalizing constant
@@
Using the approximations for the normalizing constants in this ratio, one can approximate $P(\mathbf{X}|G)$. Jones et al. (2005) and Lenkoski & Dobra (2011) both design stochastic search algorithms that in each iteration favour a move to a graph with a high posterior probability $P(G|\mathbf{X})\propto P(\mathbf{X}|G)P(G)$. However, the approximation of the normalizing constants in (12) is unstable and computationally expensive, limiting structure learning to small graphs with less than 20 nodes.
That is why others focus on inference on the joint posterior (8). For this and for inference on the precision matrix in general, it is required to sample from the $G$-Wishart distribution (10), which is challenging due to the normalizing constant (11). This led to a string of articles on efficient sampling from the $G$-Wishart distribution, see Wang & Li (2012) for a review of those. However, sampling from the $G$-Wishart distribution remained computationally expensive.
@@ 3.1 MCMC-based Approaches
Most methods in Bayesian structure learning use MCMC sampling, because of the variety of inferences it provides. Specifically, when using MCMC sampling, one can approximate the expected value of $h(G^{*})$ for any information function $h:\mathcal{G}\to\mathbb{R}$. For example, one can choose $h(G)$ equal to $1$ when an edge $e$ is in $G$, and 00 otherwise. This will give the posterior edge inclusion probabilities
@@
Similarly, with other information functions $h$, one can evaluate different posterior probabilities. For example, the probability that the true graph will have more than $10$ edges, the probability that the true graph is connected, and so on. However, this flexibility comes at a cost: MCMC algorithms are computationally demanding because they iteratively explore large parameter spaces.
@@ 3.1.1 MCMC-based approaches on the joint space
Most of the MCMC-based approaches in Bayesian structure learning create a Markov Chain $((G^{(0)},\mathbf{K}^{(0)}),\dots,(G^{(S)},\mathbf{K}^{(S)}))$ over the joint space of graphs and precision matrices that converges to the joint posterior (8). Using Equation 14 these methods can perform inference on almost all characteristics of both the graph and the precision matrix. We call these methods joint MCMC methods. They come in three types: reversible jump (RJ), birth-death (BD), and spike-and-slab (SS) methods. All joint MCMC methods are listed in Table 2 at the end of this subsection.
@@
One can define the acceptance probabilities $\alpha(G^{\prime},G)$ such that the balance condition holds and the resulting MCMC chain converges to the joint posterior. The first three RJ methods to use this strategy are the RJ-Original (RJ-O) method by Dobra et al. (2011), the RJ-Wang Li (RJ-WL) method by Wang & Li (2012), and the RJ-Cheng Lenkoski (RJ-CL) method by Cheng & Lenkoski (2012). These methods form the foundation for all later reversible jump methods. Their approach to sample from the $G$-Wishart distribution, however, is not scalable to large-scale graphs and lacks a theoretical guarantee of convergence. To address these issues, Lenkoski (2013) introduces an efficient exact sampler from the $G$-Wishart distribution. The resulting RJ-Double (RJ-D) method thereby becomes the first joint method to achieve detailed balance, and hence, theoretical convergence. Hinne et al. (2014) combine several existing techniques to improve the overall efficiency in the RJ-Double Continuous Bayes Factor (RJ-DCBF) method. Despite these improvements, RJ methods are still computationally demanding for graphs with more than 50 variables. One of the reasons is that the acceptance probabilities contain a ratio of normalizing constants that is time-consuming to evaluate. Mohammadi et al. (2023a) tackles this by developing a closed-form approximation of the ratio used in the RJ - Approximation (RJ-A) method by Mohammadi & Wit (2019). Due to the approximation, the RJ-A method loses theoretical convergence but gains computational efficiency. The RJ-Weighted Proposal (RJ-WWA) method by Van den Boom et al. (2022a) combines the best of both worlds: theoretical convergence and computational efficiency. It achieves this with two main improvements. First, the RJ-WWA algorithm is more likely to propose graphs with higher acceptance probabilities thereby improving the low acceptance probabilities of earlier algorithms. Second, they introduce a pre-acceptance step, in which the acceptance probabilities are approximated using the computationally efficient approximation of Mohammadi et al. (2023a). Only when pre-accepted, the exact acceptance probabilities are calculated. The most recent reversible jump method is the RJ-Unbiased (RJ-UB) method by Van den Boom et al. (2022b). The RJ-UB method introduces unbiased estimation to Bayesian structure learning by applying a technique of Glynn & Rhee (2014). It requires the construction of two Markov Chains $(G^{(0)},\dots,G^{(S)})$ and $(\bar{G}^{(0)},\dots,\bar{G}^{(S)})$ that meet at some time $\tau<\infty$, that is, $G^{(s)}=\bar{G}^{(s)}$ for all $s\geq\tau$. Now, for any $q=1,2,\dots$, the quantity
@@
is an unbiased estimator of $E(h(G^{*})|\mathbf{X})$. To create the two Markov chains, Van den Boom et al. (2022b) use a combination of a sequential Monte Carlo (SMC) method and a particle-independent Metropolis-Hastings method. Van den Boom et al. (2022b) are not the first to use SMC techniques in Bayesian structure learning. Tan et al. (2017) also use an SMC sampler. Their method, however, is computationally cumbersome and comes with no theoretical guarantees.
Mohammadi & Wit (2015) develop a different strategy to tackle the issue of low acceptance probabilities and create a new class of joint MCMC methods: Birth-Death (BD) methods. Their BD-Original (BD-O) method abolishes acceptance probabilities altogether while maintaining theoretical convergence. The BD-O method designs a continuous time Markov chain (Cappé et al., 2003). This chain spends a continuous time $W(s)$ in each state $(G^{(s)},K^{(s)})$ before moving to the next state. The next state is determined by adding or removing an edge based on so-called birth-death rates that are recalculated at every iteration. Its pseudo-code is shown in Algorithm 2.
@@
for state $s=1,\dots,S$ do
       Calculate the birth-death rate $r_{e}$ for every edge $e=(i,j)$;
@@
Output: Samples $(G^{(0)},\mathbf{K}^{(0)}),\dots,(G^{(S)},\mathbf{K}^{(S)})$ and waiting times $(W^{(0)},\dots,W^{(S)})$
Algorithm 2 Birth-death (BD) algorithm
The BD-O method meets the balance condition and therefore converges to the joint posterior. The desired expected value can now be obtained using
@@
where $W$ is the sum of the waiting times of all states. Per iteration, the BD-O algorithm is computationally costly because it calculates the birth and death rates for all edges. However, its MCMC chain needs less iterations to converge because it moves to a new graph every iteration. Moreover, the birth and death rates can be calculated in parallel. Just like the acceptance probabilities of the RJ methods, the birth and death rates of the BD-O method contain a ratio of normalizing constants. Using the approximation of this ratio by Mohammadi et al. (2023a), the BD-Approximation (BD-A) method achieves a significant reduction in computational cost, but the balance condition no longer holds, and there is no theoretical convergence.
Despite all the mentioned improvements, RJ and BD methods still suffer from a high computational burden. This has two reasons: first, at every iteration, a new precision matrix needs to be sampled from the $G$-Wishart distribution, which remains, despite the direct sampler of Lenkoski (2013), computationally costly. Second, at every MCMC iteration, the graph changes by at most one edge. This leads to poor mixing and convergence of the MCMC chain, especially for large-scale graphs. Both problems are overcome by a new class of joint MCMC methods called Spike-and-Slab (SS) methods. The key element of SS methods is their prior on the precision matrix, called the SS prior (Tadesse & Vannucci, 2021). An SS prior samples each element $k_{ij}$ of the precision matrix individually, such that $k_{ij}$ is likely to be close to zero for $(i,j)\not\in G$ (the spike) and likely to be away from zero for $(i,j)\in G$ (the slab). The elements of the precision matrix are still dependent, since the resulting matrix $\mathbf{K}$ needs to be positive definite. Moreover, like the $G$-Wishart prior, SS priors generally contain an intractable normalizing constant. SS priors allow for the design of efficient algorithms that sample an entirely new graph at each iteration. The Markov chain’s mixing is therefore greatly improved compared to the RJ and BD methods. Algorithm 3 presents the pseudo-code for the class of SS algorithms.
@@
for state $s=1,\dots,S$ do
       Sample a new precision matrix $\mathbf{K}^{(s)}$;
       For every edge $e=(i,j)$, calculate the probability $l_{e}$;
@@
Algorithm 3 Spike-and-slab (SS) algorithm
In the SS-Original (SS-O) method, Wang (2015) introduces SS priors to Bayesian structure learning. This prior samples off-diagonal elements $k_{ij}$ from a normal distribution with variance $v_{0}$ when $(i,j)\not\in G$ and with variance $v_{1}>v_{0}$ when $(i,j)\in G$. Diagonal elements are sampled from an exponential distribution with parameter $\eta$. The SS-Atchadé (SS-AT) method by Atchadé (2019) and the SS-BCONCORD (SS-BC) by Jalali et al. (2020) set off-diagonal elements $k_{ij}$ to zero when $(i,j)\not\in G$. These so-called discrete SS priors were introduced by Talluri et al. (2014). The SS-AT and SS-BC methods do not use the exact likelihood (9), but instead consider an approximation. These approximations are variants of the pseudo-likelihood approximation (Besag, 1975) given by
$\prod_{i=1}^{p}P(\mathbf{X}_{i}|\mathbf{X}_{-i},G,\mathbf{K})\approx P(\mathbf{X}|\mathbf{K},G),$ (17)
where $\mathbf{X}_{-i}$ denotes the data matrix $\mathbf{X}$ with the column $\mathbf{X}_{i}$ removed. Both the SS-AT and SS-BC methods motivate the use of this approximation by proving a theoretical property called posterior contraction. This means that the posterior probability of precision matrices $\mathbf{K}$ that are “far away” from the true precision matrix $\mathbf{K}^{*}$ goes to zero, as the number of observations goes to infinity. Or, put differently, the posterior distribution $P(\mathbf{K}|\mathbf{X})$ is concentrated around the true value $\mathbf{K}^{*}$. The SS-AT and SS-BC methods combine the discrete SS prior with the approximated likelihood to construct computationally efficient MCMC algorithms. On top of that, Jalali et al. (2020) prove that their SS-BC method provides a consistent estimate of the graph, i.e. as the number of observations goes to infinity, the estimated edge inclusion probabilities $p_{e}$ (13) converge (in probability) to $1$ for edges in the true graph and to 00 for edges not in the true graph.
@@
Spike-and-slab Discrete SS - O Wang (2015)
SS - AT Atchadé (2019)
SS - BC Jalali et al. (2020)*
Table 2: Overview of joint MCMC methods for Bayesian structure learning in GGMs with their corresponding prior on $\mathbf{K}$, type of Markov chain (discrete or continuous), name, and reference. The last column indicates whether an algorithm theoretically converges to the posterior distribution. Papers marked with * have not yet been published.
@@ 3.1.2 MCMC-based approaches on the graph space
A benefit of joint MCMC methods is that with the samples $(G^{(0)},\mathbf{K}^{(0)}),\dots,(G^{(S)},\mathbf{K}^{(S)})$, one can perform inference on the precision matrix. However, one is often just interested in retrieving the structure of the true graph. In that case, obtaining a new precision matrix at every iteration is a computational burden. The class of methods in this subsection overcomes this burden by creating a Markov chain $(G^{(0)},\dots,G^{(S)})$ only over the graph space. We also refer to these types of methods as G-methods.
Stingo & Marchetti (2014) introduce this strategy in their G-Stingo method. If a graph $G_{d}$ is decomposable, the pseudo-likelihood approximation (17) becomes exact. Stingo & Marchetti (2014) use this fact to design an MCMC algorithm over the space of decomposable graphs. They then design a framework to generalize this chain to the space of all graphs $\mathcal{G}$.
The pseudo-likelihood approximation (17) can also be used to approximate the marginal likelihood $P(\mathbf{X}|G)$:
@@
where $\hat{P}(\mathbf{X}|G)$ is called the marginal pseudo-likelihood (MPL). Leppä-aho et al. (2017) introduce the concept of marginal pseudo-likelihood to GGMs. They use the Wishart prior with scale parameter $b=p$ and $D=\mathbf{S}$. The Wishart prior is equivalent to the $G$-Wishart prior (10), where $G$ is the complete graph. Note that this prior uses the data $\mathbf{X}$. Such priors are called fractional priors. Leppä-aho et al. (2017) choose this prior because it leads to a closed-form expression of $\hat{P}(\mathbf{X}|G)$. They design a hill-climbing algorithm that, at each iteration, moves to a graph with a higher marginal pseudo-likelihood.
The G-MPL-Reversible Jump (G-MPLRJ) method by Mohammadi et al. (2023b) combines the closed-form expression of $\hat{P}(\mathbf{X}|G)$ with an MCMC algorithm to sample from the so-called pseudo-posterior $\hat{P}(G|\mathbf{X}):=\hat{P}(\mathbf{X}|G)P(G)/P(\mathbf{X})$. Their discrete time MCMC algorithm creates the chain $(G^{(0)},\dots,G^{(S)})$. The pseudo-code resembles Algorithm 1 with two differences. First, no precision matrix is sampled at every iteration. Second, the acceptance probabilities $\alpha(G^{\prime},G)$ are calculated using the marginal pseudo-likelihood, greatly improving the computational efficiency. The same work also presents the G-MPL-Birth Death (G-MPLBD) method. Again using the approximation in (18), this method designs a continuous time MCMC similar to Algorithm 2. Both the G-MPLRJ and the G-MPLBD algorithm meet the balance conditions and therefore converge to the pseudo-posterior. Similar to the SS - BC by Jalali et al. (2020) (see Subsection 3.1.1), Mohammadi et al. (2023b) prove two theoretical results: the pseudo-posterior concentrates around the true graph $G^{*}$ and the edge inclusion probabilities are consistent.